# 2.5cm2.5cm2cm2.0cm

### v1

@@ Abstract
Femtoscopy is a unique tool to investigate the space-time geometry of the matter created in ultra-relativistic collisions. If the probability density distribution of hadron emission is parametrized, then the dependence of its parameters on particle momentum, collision energy, and collision geometry can be given. In recent years, several measurements came to light that indicated the adequacy of assuming a Lévy-stable shape for the mentioned distribution. In parallel, several new phenomenological developments appeared, aiding the interpretation of the experimental results, or providing tools for the measurements. In this paper we review and discuss some of these advances, phenomenological and experimental.
@@ 1 Introduction
In high-energy physics, femtoscopy [1, 2] is the most important way to gain information about the space-time geometry of the particle-emitting source (the probability density of particle creation in an infinitesimal interval in phase-space). Historically, the phenomenon on which this technique is based was discovered by R. Hanbury Brown and R. Q. Twiss in astronomy in two-photon correlations [3]. Hence measurements in this field are often labeled “HBT-correlations”. Independently, proton-proton collisions were investigated by G. Goldhaber and collaborators, and correlations of identical pions were observed [4]. Our current understanding suggests that both phenomena are based on the wave-like nature of light and pions, or alternatively (starting from the quantum nature of particle detection) on the Bose-Einstein symmetrization of pair wavefunctions. See, for example, the introduction of Refs. [5, 6] for details.
Two-particle momentum correlations (or Bose-Einstein correlations) are defined as
@@
where $p_{1}$ and $p_{2}$ are the momenta of the two particles, $N_{12}$ is the invariant pair-momentum distribution, while $N_{1}$ and $N_{2}$ are the invariant momentum distributions of the particles $1$ and $2$ (for identical particles, these distributions are of course also identical). In general, there are several reasons for correlated particle production (i.e., $C_{2}\neq 1$): for example collective flow, jets, resonance decay, and conservation laws. In the case of high-multiplicity heavy-ion collisions, the main source of correlations is the above-mentioned HBT effect, and they appear mostly at small relative momenta, where $p_{1}\approx p_{2}$, or if one introduces $q=p_{1}-p_{2}$ and $K=(p_{1}-p_{2})/2$, the criterion $q\ll K$ may be given. It turns out that in the absence of final-state interactions and multiparticle-correlations, a simple relation can be established [7, 8, 9] between the measured momentum correlation $C_{2}(q,K)$ and the spatial pair-distribution (or pair source) $D(r,K)$ (indicating the probability density of creating a pair in an infinitesimal interval around distance $r$ and mean momentum $K$), given as
@@
where the $\widetilde{D}$ denotes the Fourier-transform of the pair source $D(r,K)$ in its first variable (see a comprehensive overview of this result in Ref. [8]. It is characteristic to sources in heavy ion collisions that the dependence of $\widetilde{D}(q,K)$ and hence of $C_{2}(q,K)$ on $K$ is smoother than on $q$, this property is often called the “smoothness approximation” [2]. One may furthermore note that the strength of the correlation functions may be decreased by several phenomena: decays of long-lived resonances [10], coherence [11], quasi-random electromagnetic fields [12]. Taking these into account the above equation is modified as
@@
where $\lambda$ is the correlation strength parameter, in the core-halo model [10] identified with the squared fraction of primordial pions amongst all produced pions.
If one aims at characterizing the geometry of the (pair) source, one often assumes a parametrization of it, in such a way that the source shape is a parametric distribution as a function of $r$, and its parameters depend on the pair momentum $K$. A usual assumption (see e.g. [13]) is that of a Gaussian source, in which case
@@
where $R(K)$ is the momentum-dependent width of the source (here a factor of $2$ would appear in the denominator for the single-particle source, which then cancels from the pair source), which can also be called the HBT or femtoscopic size. This would lead to a correlation function of
@@
This HBT radius is of high relevance in terms of the physics of the quark-gluon plasma, as (especially if measured in multiple dimensions) it sheds light on transverse expansion or details of the phase diagram, as indicated, e.g., in Refs. [14, 15, 16, 17, 18]. We note here that usually instead of pair momentum $K$, pair transverse mass $m_{T}=\sqrt{K_{T}^{2}+m^{2}}$ is defined (in $c=1$ units), where $K_{T}$ is the transverse component of $K$, and $m$ is the mass of the particle being investigated.
Note finally that the above is true for non-interacting particles. However, even pions undergo electromagnetic final-state interactions. The effect of these is usually handled via the so-called Coulomb-correction [8]; while the strong interaction of pions is usually negligible [19], but is much more important for baryons and in case of non-identical correlations [20, 21].
@@ 2 Basics of femtoscopy with Lévy sources
As mentioned above, a usual assumption for the shape of the (pair or single-particle) source is the Gaussian distribution. However, the observation of a long tail on the reconstructed source shape suggested [22, 23] that for an adequate description, one can and should go beyond the Gaussian approximation. A plausible generalization is to utilize Lévy-stable distributions, as introduced in Ref. [24] and utilized, for example, in Ref. [6]. These appear when the conditions of the Central Limit Theorem are not met, in particular when the independent and identically distributed random variables (the sum of which we investigate) do not have a finite mean or variance. In this case, the Generalized Central Limit Theorem applies [25], and the limiting distributions are the so-called Lévy-stable distributions [26]. The spherically symmetric Lévy-stable distributions (or Lévy distributions for short) are defined as
@@
where $\mathbf{r}$ is the spatial vector, the variable of the distribution, and $\alpha$ is the Lévy exponent, characterizing the shape of the distribution; the stability criterion (i.e., the limiting nature of the distribution) is met if $0<\alpha\leq 2$. Furthermore, $R$ is the spatial scale of the distribution, characterizing its size (even if the variance or root-mean-square is infinity), and corresponding to the HBT or femtoscopic radius in our case. Let us note here that if $\alpha<2$, then the distribution exhibits a power-law tail in $r=|\mathbf{r}|$, with $\mathcal{L}(\alpha,R,\mathbf{r})\propto(r/R)^{-(3+\alpha)}$. Its angle-averaged version behaves as $r^{2}\mathcal{L}(\alpha,R,\mathbf{r})\propto r^{-1-\alpha}$; underlining the fact that these distributions have no finite variance for $\alpha<2$, and $r$ has no finite mean for $\alpha<1$ (clearly the mean of $\mathbf{r}$ is zero, due to spherical symmetry).
Figure 1: The left plot shows correlation functions $C(Q)$ (equivalent to $C_{2}(q)$ as denoted in the text) based on various source geometries: a 3D angle averaged Gaussian with $R_{\rm out}=6$ fm, $R_{\rm side}=4$ fm, and $R_{\rm long}=8$ fm, a spherically symmetric (‘1D’) Gaussian with $R=6$ fm, and an also spherically symmetric (‘1D’) Lévy with $\alpha=1.4$ and $R=6$ fm. The right plot shows the ratio of the correlation functions based on a 3D Gaussian source and the spherically symmetric Lévy source, divided by the spherically symmetric Gaussian source.
The appearance of Lévy distributions in high-energy physics may be caused by several (competing) physical phenomena:
  • Anomalous diffusion [27, 28], or Lévy-flight, probably due to hadronic rescattering; if this happens in the hadronic stage then an ordering of the Lévy exponent is expected as $\alpha({\rm kaons})<\alpha({\rm pions})<\alpha({\rm protons})$, due to the cross-section and hence mean free path of the given particle species. [28]
@@
  • Second-order phase transitions [31], where the correlation length diverges near the critical point, and at the critical point the spatial correlations exhibit a power-law tail with exponent $\eta$; this is one of the critical exponents, and its value is suggested to be $0{,}03631(3)$ [32] for the 3D Ising model, or $0{,}50\pm 0{,}05$ for the random external field 3D Ising model [33] – QCD is expected to be in the same universality class as one of them [34, 35].
@@
  • Resonance decays [28, 36, 9], where the power-law tail is generated by the set of resonances decaying into pions (or the given investigated particle species); this phenomenon is similar to the decay heat of used fuel rods in power plants, see Fig. 1. of Ref. [37]. Note that the simulations of Refs [9, 36] indicate Lévy distributions in EPOS even before resonance decays, hence (at least in EPOS) these cannot be the only reason for the appearance of Lévy distributions.
@@
Further ideas have been promoted recently, for example, event averaging and angular averaging, see Ref. [38, 39]. Note that in Refs. [36, 9] source shapes from the EPOS model have been analyzed on an event-by-event basis, and Lévy distributions appeared in individual events as well. Note furthermore that while a non-spherical Gaussian source is clearly exhibiting a non-Gaussian shape when measured in an angle-averaged fashion, a spherically symmetric Lévy distribution leads to quite different correlation functions, as illustrated in Fig. 1. This shows that a 3D angle-averaged Gaussian (with radii $R_{\rm out}=6$ fm, $R_{\rm side}=4$ fm, and $R_{\rm long}=8$ fm in the Bertsch-Pratt frame [40, 41]) source leads to a non-Gaussian 1D correlation function, however, this apparent “non-Gaussianity” is distinctly different from a 1D (or spherically symmetric) correlation function based on a Lévy distribution. With respect to this, it shall also be noted that 3D HBT measurements in the longitudinally comoving system (LCMS) have been performed in PHENIX [42] and the resulting $\alpha$ exponent appeared to be the same as the one from the 1D analysis at [6], indicating that it is not non-sphericality that causes the appearance of Lévy distributions. Note that a further difference between 1D and 3D analyses is, that in 1D, one cannot easily make the transformation from the LCMS to the pair rest frame, where the Coulomb correction is calculated (the two-particle Schrödinger equation is a good approximation to the given problem only in this frame, where the nonrelativistic formulas may be utilized). It was, however, found in Ref. [7] that simple but adequate approximation formulas can be found, in particular for the source radius to be used in the Coulomb correction calculation.
@@ 3 Measures of the source extent
[…]
The source, as discussed above, may exhibit various shapes, and even in the absence of a finite variance, one is interested in finding an “assumption-independent” measure of the source extent. One of these may be the half with at half maximum (HWHM), the distance (from the origin or center) where the probability density decreases to half of the maximal value (admitted at the origin or center usually). Another such measure may be the half width at half-integral, i.e., the radius of the region from where half of the particles are emitted. In the case of a Gaussian, the HWHM of the single-particle source can be readily calculated as
@@
The above results can be understood for a Lévy distribution with $\alpha=2$. However, these values strongly depend on $\alpha$, as illustrated by Fig. 2. This means that either HWHI or HWHM can change
  • due to a change in size (i.e., the $R_{\textrm{Lévy}}$ scale), or
  • due to a change in shape.
@@ 4 Signs of the critical point in femtoscopy
As noted above, the Lévy exponent $\alpha$ was identified in Ref. [31] with the critical exponent of spatial correlations, usually denoted by $\eta$. It was furthermore suggested that while far from the critical point, $\alpha$ may be of a value close to 2, closer to the critical point, it may decrease to the above-mentioned values, close to or below 0.5. This behavior may be blurred by finite size effects; however, Refs. [44, 45] found that the correlation exponent $\eta$ is largely unchanged by finite size effects in the 3D Ising model and also in the random field 3D Ising model. This suggests that it is not a priori impossible to search for the critical point via the measurement of $\alpha$. Furthermore, discussed, e.g., in Ref. [14], femtoscopic radii may be interpreted as signaling the vicinity of the critical point. Finally, one may note that the HBT radii are also connected to the lifetime of the quark medium [46], which is also influenced by the equation of state of the medium. Hence, the collision energy dependence of HBT radii serves as an imporant tool to investigate the QCD phase diagram.
Figure 3: Varouios source geometry characteristics versus collision energy ($\sqrt{s_{{}_{NN}}}$) in a hypothetical scenario where $\alpha$ (red crosses, right vertical axis) has a minimum and $R_{\rm Levy}$ a maximum near $\sqrt{s_{{}_{NN}}}=20$ GeV. The following calculated source measures were calculated and are shown: half width at half integral ($R_{\rm HWHI}$, orange downward-facing triangles), half width at half maximum ($R_{\rm HWHM}$, green upward-facing triangles), and a radius from a forced Gaussian fit ($R_{\rm Gaussfit}$, purple squares).
Let us now suppose a scenario where the Lévy scale $R_{\textrm{Lévy}}$ exhibits a maximum near $\sqrt{s_{{}_{NN}}}=20$ GeV, while the Lévy exponent $\alpha$ has a minimum at the same location, as shown in Fig. 3. Let us then generate correlation functions with a toy model where the supposed Lévy parameters are generated with realistic statistical uncertainties (following those from experimental data, such as the one in Ref. [6]). We investigate three different size measures (besides $R_{\textrm{Lévy}}$, which is an input, but it was confirmed that from a Lévy fit it can be retained):
@@
Fig. 2 shows the results, and the observation is that $R_{\rm Gauss\;fit}$, contrary to the input $R_{\textrm{Lévy}}$, shows an approximately monotonic increase with $\sqrt{s_{{}_{NN}}}$. On the other hand, there is a minimum in HWHM, and HWHI exhibits a trend change.

### v3

@@ Abstract
Femtoscopy is a unique tool to investigate the space-time geometry of the matter created in ultra-relativistic collisions. If the probability density distribution of hadron emission is parametrized, then the dependence of its parameters on particle momentum, collision energy, and collision geometry can be given. In recent years, several measurements came to light that indicated the adequacy of assuming a Lévy-stable shape for the mentioned distribution. In parallel, several new phenomenological developments appeared, aiding the interpretation of the experimental results or providing tools for the measurements. In this paper, we discuss important aspects of femtoscopy with Lévy sources in light of some of these advances, including phenomenological and experimental ones.
@@ 1 Introduction
In high-energy physics, femtoscopy [1, 2] is the most important way to gain information about the space-time geometry of the particle-emitting source (the probability density of particle creation in an infinitesimal interval in phase-space). Historically, the phenomenon on which this technique is based was discovered by R. Hanbury Brown and R. Q. Twiss in astronomy in two-photon correlations [3]. Hence, measurements in this field are often labeled “HBT-correlations”. Independently, G. Goldhaber and collaborators investigated proton–proton collisions and observed the correlations of identical pions [4]. Our current understanding suggests that both phenomena are based on the wave-like nature of light and pions, or alternatively (starting from the quantum nature of particle detection) on the Bose–Einstein symmetrization of pair wavefunctions. See, for example, the introduction of Refs. [5, 6] for details.
Two-particle momentum correlations (or Bose–Einstein correlations) are defined as
@@
where $p_{1}$ and $p_{2}$ are the momenta of the two particles, $N_{12}$ is the invariant pair-momentum distribution, while $N_{1}$ and $N_{2}$ are the invariant momentum distributions of the particles $1$ and $2$ (for identical particles, these distributions are of course also identical). In general, there are several reasons for correlated particle production (i.e., $C_{2}\neq 1$): for example, collective flow, jets, resonance decay, and conservation laws. In the case of high-multiplicity heavy-ion collisions, the main source of correlations is the HBT effect mentioned above, and they appear mostly at small relative momenta, where $p_{1}\approx p_{2}$, or if one introduces $K=(p_{1}+p_{2})/2$ and $q=p_{1}-p_{2}$, the criterion $q\ll K$ may be given. It turns out that in the absence of final-state interactions and multiparticle-correlations, a simple relation can be established [7, 8, 9] between the measured momentum correlation $C_{2}(q,K)$ and the spatial pair-distribution (or pair source) $D(r,K)$ (indicating the probability density of creating a pair in an infinitesimal interval around distance $r$ and mean momentum $K$), given as
@@
where $\widetilde{D}$ denotes the Fourier-transform of the pair source $D(r,K)$ in its first variable (see a comprehensive overview of this result in Ref. [8]). It is characteristic to sources in heavy-ion collisions that the dependence of $\widetilde{D}(q,K)$ and, hence, of $C_{2}(q,K)$ on $K$ is smoother than on $q$; this property is often called the “smoothness approximation” [2]. One may furthermore note that several phenomena may decrease the strength of the correlation functions: decays of long-lived resonances [10], coherence [11], and quasi-random electromagnetic fields [12]. Taking these into account, the above equation is modified as
@@
where $\lambda$ is the correlation strength parameter in the core–halo model [10] identified with the squared fraction of primordial pions amongst all produced pions.
Suppose one aims to characterize the geometry of the (pair) source. In that case, one often assumes a parametrization of it, in such a way that the source shape is a parametric distribution as a function of $r$, and its parameters depend on the pair momentum $K$. A usual assumption (see e.g., [13]) is that of a Gaussian source, in which case
@@
where $R(K)$ is the momentum-dependent width of the source (here, a factor of $2$ would appear in the denominator for the single-particle source, which then cancels from the pair source), which can also be called the HBT or femtoscopic size. This would lead to a correlation function of
@@
This HBT radius is of high relevance in terms of the physics of the quark–gluon plasma, as (especially if measured in multiple dimensions) it sheds light on transverse expansion or details of the phase diagram, as indicated, e.g., in Refs. [14, 15, 16, 17, 18]. We note here that, usually instead of pair momentum $K$, pair transverse mass $m_{T}=\sqrt{K_{T}^{2}+m^{2}}$ is defined (in $c=1$ units), where $K_{T}$ is the transverse component of $K$, and $m$ is the mass of the particle being investigated.
Note finally that the above is true for non-interacting particles. However, even pions undergo electromagnetic final-state interactions. The effect of these is usually handled via the so-called Coulomb-correction [8]. At the same time, the strong interaction of pions is usually negligible [19], but is much more important for baryons and in case of non-identical correlations [20, 21].
@@ 2 Basics of Femtoscopy with Lévy Sources
As mentioned above, a usual assumption for the shape of the (pair or single-particle) source is the Gaussian distribution. However, the observation of a long tail on the reconstructed source shape suggested [22, 23] that for an adequate description, one can and should go beyond the Gaussian approximation. A plausible generalization is to utilize Lévy-stable distributions, as introduced in Ref. [24] and utilized, for example, in Ref. [6]. These appear when the conditions of the Central Limit Theorem are not met, particularly when the independent and identically distributed random variables (the sum of which we investigate) do not have a finite mean or variance. In this case, the Generalized Central Limit Theorem applies [25], and the limiting distributions are the so-called Lévy-stable distributions [26]. The spherically symmetric Lévy-stable distributions (or Lévy distributions for short) are defined as
@@
where $\mathbf{r}$ is the spatial vector, the variable of the distribution, and $\alpha$ is the Lévy exponent, characterizing the shape of the distribution; the stability criterion (i.e., the limiting nature of the distribution) is met if $0<\alpha\leq 2$. Furthermore, $R$ is the spatial scale of the distribution, characterizing its size (even if the variance or root-mean-square is infinity) and corresponding to the HBT or femtoscopic radius in our case. Let us note here that if $\alpha<2$, then the distribution exhibits a power-law tail in $r=|\mathbf{r}|$, with $\mathcal{L}(\alpha,R,\mathbf{r})\propto(r/R)^{-(3+\alpha)}$. Its angle-averaged version behaves as $r^{2}\mathcal{L}(\alpha,R,\mathbf{r})\propto r^{-1-\alpha}$, underlining the fact that these distributions have no finite variance for $\alpha<2$, and $r$ has no finite mean for $\alpha<1$ (clearly, the mean of $\mathbf{r}$ is zero, due to spherical symmetry).
Several (competing) physical phenomena may cause the appearance of Lévy distributions in high-energy physics:
  • Anomalous diffusion [27, 28], or Lévy-flight, probably due to hadronic rescattering; if this happens in the hadronic stage, then an ordering of the Lévy exponent is expected as $\alpha({\rm kaons})<\alpha({\rm pions})<\alpha({\rm protons})$, due to the cross-section and, hence, mean free path of the given particle species. [28]
@@
  • Second-order phase transitions [31], where the correlation length diverges near the critical point, and at the critical point, the spatial correlations exhibit a power-law tail with exponent $\eta$; this is one of the critical exponents, and its value is suggested to be $0.03631(3)$ [32] for the 3D Ising model, or $0.50\pm 0.05$ for the random external field 3D Ising model [33]—QCD is expected to be in the same universality class as one of them [34, 35].
@@
  • Resonance decays [28, 36, 9], where the power-law tail is generated by the set of resonances decaying into pions (or the given investigated particle species); this phenomenon is similar to the decay heat of used fuel rods in power plants, see Figure 1 of Ref. [37]. Note that the simulations of Refs [9, 36] indicate Lévy distributions in EPOS even before resonance decays; hence (at least in EPOS), these cannot be the only reason for the appearance of Lévy distributions.
@@
Further ideas have been promoted recently, for example, event averaging and angular averaging, see Ref. [38, 39]. Note that in Refs. [36, 9], source shapes from the EPOS model have been analyzed on an event-by-event basis, and Lévy distributions appeared in individual events as well. Note furthermore that while a non-spherical Gaussian source is clearly exhibiting a non-Gaussian shape when measured in an angle-averaged fashion, a spherically symmetric Lévy distribution leads to quite different correlation functions, as illustrated in Figure 1. This shows that a 3D angle-averaged Gaussian (with radii $R_{\rm out}=6$ fm, $R_{\rm side}=4$ fm, and $R_{\rm long}=8$ fm in the Bertsch–Pratt frame [40, 41]) source leads to a non-Gaussian 1D correlation function. However, this apparent “non-Gaussianity” is distinctly different from a 1D (or spherically symmetric) correlation function based on a Lévy distribution. With respect to this, it shall also be noted that 3D HBT measurements in the longitudinally comoving system (LCMS) have been performed in PHENIX [42] and the resulting $\alpha$ exponent appeared to be the same as the one from the 1D analysis at [6], indicating that it is not non-sphericality that causes the appearance of Lévy distributions. Note that a further difference between 1D and 3D analyses is that, in 1D, one cannot easily make the transformation from the LCMS to the pair rest frame, where the Coulomb correction is calculated (the two-particle Schrödinger equation is a good approximation to the given problem only in this frame, where the nonrelativistic formulas may be utilized). However, in Ref. [7], simple but adequate approximation formulas were found, particularly for the source radius to be used in the Coulomb correction calculation.
Figure 1: The left plot shows correlation functions $C(Q)$ (equivalent to $C_{2}(q)$ as denoted in the text) based on various source geometries: a 3D angle-averaged Gaussian with $R_{\rm out}=6$ fm, $R_{\rm side}=4$ fm, and $R_{\rm long}=8$ fm, a spherically symmetric (‘1D’) Gaussian with $R=6$ fm, and an also spherically symmetric (‘1D’) Lévy with $\alpha=1.4$ and $R=6$ fm. The right plot shows the ratio of the correlation functions based on a 3D Gaussian source and the spherically symmetric Lévy source, divided by the spherically symmetric Gaussian source.
@@ 3 Measures of the Source Extent
As discussed above, the source may exhibit various shapes, and even in the absence of a finite variance, one is interested in finding an “assumption-independent” measure of the source extent. One of these may be the half with at half maximum (HWHM), the distance (from the origin or center) where the probability density decreases to half of the maximal value (admitted at the origin or center usually). Another such measure may be the half width at half-integral, i.e., the radius of the region from where half of the particles are emitted. In the case of a Gaussian, the HWHM of the single-particle source can be readily calculated as
@@
The above results can be understood for a Lévy distribution with $\alpha=2$. However, these values strongly depend on $\alpha$, as illustrated by the left plot of Figure 2. This means that either HWHI or HWHM can change due to either of the following:
  • A change in size (i.e., the $R_{\textrm{Lévy}}$ scale);
  • A change in shape.
@@
The right plot of Figure 2 furthermore indicates the ratio of Lévy and Gaussian radii, when the HWHI or the HWMH of the two distributions is the same. This indicates that some sort of “translation” can be found between Lévy (free $\alpha$) and Gaussian ($\alpha=2$) radii: for example at $\alpha=1.3$, the Gaussian distribution with the same HWHM is of radius $R_{\rm 3D\;Gauss}\approx 1.94R_{\alpha=1.3}$, while it has the same HWHI if $R_{\rm 3D\;Gauss}\approx 1.21R_{\alpha=1.3}$.
[…]
@@ 4 Signs of the Critical Point in Femtoscopy
As noted above, the Lévy exponent $\alpha$ was identified in Ref. [31] with the critical exponent of spatial correlations, usually denoted by $\eta$. It was furthermore suggested that, while far from the critical point, $\alpha$ may be of a value close to 2, closer to the critical point, or it may decrease to the above-mentioned values, close to or below 0.5. This behavior may be blurred by finite size effects; however, Refs. [44, 45] found that the correlation exponent $\eta$ is largely unchanged by finite size effects in the 3D Ising model and also in the random field 3D Ising model. This suggests that it is not a priori impossible to search for the critical point via the measurement of $\alpha$. Furthermore, discussed, e.g., in Ref. [14], femtoscopic radii may be interpreted as signaling the vicinity of the critical point. Finally, one may note that the HBT radii are also connected to the lifetime of the quark medium [46], which is also influenced by the equation of state of the medium. Hence, the collision energy dependence of HBT radii serves as an important tool to investigate the QCD phase diagram.
Let us now suppose a scenario where the Lévy scale $R_{\textrm{Lévy}}$ exhibits a maximum near a collision energy of $\sqrt{s_{{}_{NN}}}=20$ GeV, while the Lévy exponent $\alpha$ has a minimum at the same location, as shown in Figure 3. The values depicted on this plot are purely hypothetical, assumed to take the value shown on the figure for illustration purposes. Let us then generate correlation functions with a toy model where the supposed Lévy parameters are generated with realistic statistical uncertainties (following those from experimental data, such as the one in Ref. [6]). We investigate three different size measures (besides $R_{\textrm{Lévy}}$, which is an input, but it was confirmed that from a Lévy fit it can be retained):
@@
Figure 3 shows the results, and the observation is that $R_{\rm Gauss\;fit}$, contrary to the input $R_{\textrm{Lévy}}$, shows an approximately monotonic increase with $\sqrt{s_{{}_{NN}}}$. On the other hand, there is a minimum in HWHM, and HWHI exhibits a trend change.
@@
Figure 3: Various source geometry characteristics versus collision energy ($\sqrt{s_{{}_{NN}}}$) in a hypothetical scenario where $\alpha$ (red crosses, right vertical axis) has a minimum and $R_{\rm Levy}$ a maximum near $\sqrt{s_{{}_{NN}}}=20$ GeV. The following calculated source measures were calculated and are shown: half width at half integral ($R_{\rm HWHI}$, orange downward-facing triangles), half width at half maximum ($R_{\rm HWHM}$, green upward-facing triangles), and a radius from a forced Gaussian fit ($R_{\rm Gaussfit}$, purple squares).